Use big š‘€ method to solve Minimize š‘ = 6š‘„1 + 3š‘„2 + 4š‘„3 š‘†š‘¢š‘š‘—š‘’š‘š‘” š‘”š‘œ



š‘„1 ā‰„ 30; š‘„2 ā‰„ 50; š‘„3 ā‰„ 20; š‘„1 + š‘„2 + š‘„3 = 120

1
Expert's answer
2021-11-29T16:56:36-0500

Minimize "\ud835\udc4d = 6\ud835\udc65_1 + 3\ud835\udc65_2 + 4\ud835\udc65_3"

š‘†š‘¢š‘š‘—š‘’š‘š‘” š‘”š‘œ

"\ud835\udc651 \u2265 30"

"\ud835\udc652 \u2265 50"

"\ud835\udc653 \u2265 20"

"\ud835\udc65_1 + \ud835\udc65_2 + \ud835\udc65_3 = 120"


After introducing surplus,artificial variables

MaxĀ Z=6x1+3x2+4x3+0S1+0S2+0S3-MA1-MA2-MA3-MA4

subject to x1-S1+A1=30

x2-S2+A2=50

x3-S3+A3=20

x1+x2+x3+A4=120

andĀ x1,x2,x3,S1,S2,S3,A1,A2,A3,A4ā‰„0



Negative minimumĀ Zj-CjĀ isĀ -2M-6Ā and its column index isĀ 1. So,Ā the entering variable isĀ x1.

Minimum ratio isĀ 30Ā and its row index isĀ 1. So,Ā the leaving basis variable isĀ A1.

āˆ“Ā The pivot element isĀ 1.

EnteringĀ =x1, DepartingĀ =A1, Key ElementĀ =1



Negative minimumĀ Zj-CjĀ isĀ -2M-4Ā and its column index isĀ 3. So,Ā the entering variable isĀ x3.

Minimum ratio isĀ 20Ā and its row index isĀ 3. So,Ā the leaving basis variable isĀ A3.

āˆ“Ā The pivot element isĀ 1.

EnteringĀ =x3, DepartingĀ =A3, Key ElementĀ =1



Negative minimumĀ Zj-CjĀ isĀ -2M-3Ā and its column index isĀ 2. So,Ā the entering variable isĀ x2.

Minimum ratio isĀ 50Ā and its row index isĀ 2. So,Ā the leaving basis variable isĀ A2.

āˆ“Ā The pivot element isĀ 1.

EnteringĀ =x2, DepartingĀ =A2, Key ElementĀ =1



Negative minimumĀ Zj-CjĀ isĀ -M-6Ā and its column index isĀ 4. So,Ā the entering variable isĀ S1.

Minimum ratio isĀ 20Ā and its row index isĀ 4. So,Ā the leaving basis variable isĀ A4.

āˆ“Ā The pivot element isĀ 1.

EnteringĀ =S1, DepartingĀ =A4, Key ElementĀ =1



Since allĀ Zj-Cjā‰„0

Hence, optimal solution is arrived with value of variables as :

x1=50,x2=50,x3=20

MaxĀ Z=530